what is.....the answer to: How much oil should be mixed with 20 gallons of gas if the recommened ratio of gasoline and oil is 50 to 1?
Answer by rahk(1)

An easy way of solving this problem is to use a proportion. Here is how I did it.

50/1 = 20/x

20 x 1 = 20
20/50 = 40/100 = .4
x = .4
Hope that helps.